two consecutive odd integers have a product of 99. find all sets of integers that satisfy this decription
Answer by stanbon(75887)

two consecutive odd integers have a product of 99. find all sets of integers that satisfy this decription
-----
1st: 2x-1
2nd: 2x+1
-----
Equation:
(2x-1)(2x+1) = 99
4x^2-1 = 99
4x^2 = 100
x^2 = 25
x = 5 or x = -5
====================
If x = 5,
2x-1 = 9
2x+1 = 11
---
If x = -5
2x-1 = -11
2x+1 = -9
=============
Cheers,
Stan H.
==============